05 March 2014

Last day at PEARL with the full Intensive Phase team

On route to the Ridge Lab this morning, the team had their first wildlife sighting: a ptarmigan along the side of the road. Ptarmigan are the official bird of Nunavut. Alas, it didn't bring the FTS team much luck. The weather has continued to pose challenges. We had another cloudy day. Eureka and PEARL's temperature was -30°C today, with light snow. This isn't too bad, although winds caused tonight's ACE ozonesonde launch to fail.

In the PEARL Ridge Lab's IR lab, Joseph and Dan took 7 MIR measurements with the Bruker during a brief window of time when the Sun broke through the clouds. Dan was able to take 15 measurements with PARIS. With Paul's help, Dan moved the old suntracker equipment into a storage cabinet. This makes it easier for us to move around a critical area of the IR lab.

Over in the UV-Vis Lab, Sophie discovered that the new diffuser she put on the UT-GBS telescope decreases the light signal. This was expected. However, analysis of the tests she did showed that there might be a light leak in the spectrometer. Sophie covered the spectrometer with black cloth and used a flash light to identify the leak. There is a very narrow space between the lid and the spectrometer that allows stray light to enter. Sophie has added black tape and black cloth to fully cover the spectrometer and avoid this problem. New tests will be done to assess the quality of the measurements tomorrow.

Sophie and Paul also went on the roof to further secure SAOZ.

Zahra finally managed to align the SPS and MAESTRO tracker with the sun. All instruments are working nominally and taking measurements. The software will run automatically for the rest of the campaign. Since Zahra and Tom are leaving tomorrow, other team members will check on the computers and instruments daily. The instrument apertures must have frost and ice cleaned off of them. If there is a power failure or software freezes, the instruments and tracker and their software will need to be restarted.

The Brewer is fully operational but is waiting for clear skies to do a sun-pointing alignment (sun sighting). At the moment, it is approximately pointing at the sun.

Down at 0PAL, the CRL took 17.5 hours of data in all UV and Vis channels. After a complete shutdown of the lidar, Sham and Emily went through the optics to check whether there is any damage or dust inside. Sham also reviewed how to remove the pump chamber in the UV laser. She is ready to take the CRL reins for the rest of the campaign!

A produce flight landed at Eureka today, bringing fresh supplies and a ride home for half our team. Tomorrow morning, Tom, Zahra, and Emily will be departing, along with some of the Eureka Weather Station staff. The rest of the intensive phase team (Dan, Joseph, Paul, Sophie, and Sham) will remain until their flight out of Eureka on March 19.

- Dan
Back to Daily Reports
Back to Home